\begin{thebibliography}{mm}

\bibitem{RY_99} D. Revuz and M. Yor, ``Continuous martingales and
Brownian motion'', Grundlehren der mathematischen Wissenschaften 293,
Springer, 1999.

\bibitem{Ru_13}J.  Ruf, {\em A new proof for the conditions of
Novikov and Kazamaki\/}, Stochastic Process. Appl., Vol 123 (2013),
No. 2, 404--421.

\bibitem{Ru_15} J. Ruf, {\em The martingale property in the context
of stochastic differential equations\/}, Electron. Commun. Probab.,
Vol. 20
(2015), No. 34, 10 pp.

\bibitem{Ka_77} N. Kazamaki   {\em
On a problem of Girsanov\/}, T\^ohoku Math. Journ.,
 Vol. 29 (1977), 597-600.

\bibitem{Kr_95} N.V. Krylov,
``Introduction to the theory of diffusion processes'', Amer.
Math. Soc., Providence, RI, 1995.
 

\bibitem{LSh_72} R.Sh. Liptser and A.N. Shiryaev,
{\em
 The absolute continuity with respect to 
Wiener measure of measures that correspond 
to processes of diffusion type\/}, Izv. Akad. Nauk SSSR 
Ser. Mat., Vol. 36 (1972), 847--889 in  Russian; English translation
in Math. USSR-Izv., Vol. 6 (1972), 839--882.

\bibitem{LSh_01} R.Sh. Liptser and A.N. Shiryaev,
``Statistics of random processes. I. General theory'',
 Translated from
the 1974 Russian original by A. B. Aries. Second, revised and
expanded edition. Applications of Mathematics (New York), 5.
Stochastic Modelling and Applied Probability. Springer-Verlag,
Berlin, 2001.



\bibitem{No_72} A.A. Novikov,  
{\em  A certain identity for stochastic integrals\/}, Teor. 
Verojatnost. i Primenen., Vol. 17 (1972), 761--765
in Russian; English translation in
 Theory Probab. Appl., Vol. 17 (1973),
No. 4, 717--720.

\bibitem{No_79} A.A. Novikov,  
{\em 
Conditions for uniform integrability of continuous
 nonnegative martingales\/}, Teor. Veroyatnost. i Primenen.,
Vol. 24 (1979), No. 4, 821--825 in Russian; English translation in
  Theory Probab. Appl., Vol. 24 (1979), No. 4, 824--829.
\end{thebibliography}

\documentclass[11pt,reqno]{amsart}

\newtheorem{theorem}{Theorem}[section]

\theoremstyle{definition}

\theoremstyle{remark}

\theoremstyle{remark}
\newtheorem{example}[theorem]{Example}

\makeatletter
\def\dashint{\operatorname%
{\,\,\text{\bf--}\kern-.98em\DOTSI\intop\ilimits@\!\!}}
\makeatother

\newcommand{\nliminf}{\operatornamewithlimits{\underline{lim}}}

\newcommand\cF{\mathcal{F}}

 \newcommand{\mysection}[1]{\section{#1}
 \setcounter{equation}{0}}

\title[Result of A. Novikov
and Girsanov's theorem]{A few comments on a result of A. Novikov
and Girsanov's theorem}
\author[]{N.V. Krylov} 
\address{127 Vincent Hall, University of Minnesota,
 Minneapolis, MN, 55455}
 
\email{nkrylov@umn.edu}
\keywords{Exponential martingales, Novikov's condition, Girsanov's
theorem}
\date{}%

\subjclass{60G44, 60H10}

\begin{document}

\begin{abstract}
 
We give a simple proof that for a continuous 
local martingale $M_{t}$ 
$$
\nliminf_{\varepsilon\downarrow0}\varepsilon
\log Ee^{(1-\varepsilon)
\langle M\rangle_{\infty}/2}<\infty
\Longrightarrow E\exp(M_{\infty}-\langle
 M\rangle_{\infty}/2)=1.
$$
 
\end{abstract}

\maketitle

\mysection{Main Result}

  Let $(\Omega,\mathcal{F},P)$ be a complete 
probability space and let
$M_{t}$ be a continuous local martingale 
on $(\Omega,\mathcal{F},P)$, provided with an appropriate
filtration of sub $\sigma$-fields of $\cF$,
such that $\langle M\rangle=\langle M
\rangle_{\infty}<\infty$ (a.s.).
Define
$$
M=M_{\infty},\,\,\,\rho=\rho(M)=e^{M-
\langle M\rangle/2},\,\,\,\rho_{t}=\rho_{t}
(M)=e^{M_{t}-
\langle M\rangle_{t}/2}.
$$
The process $\rho_{t}$ is called an exponential martingale,
although it is not necessarily a martingale.
We will be discussing generalizations of the following 
celebrated
 result of A.~Novikov (1972), which gives a sufficient
condition for $\rho_{t}$  to be a martingale:
\begin{equation}
                                           \label{10.17.1}
Ee^{\langle M\rangle/2}<\infty\Longrightarrow 
E\rho=1.
\end{equation}
This result is quite important in many applications
related to absolute continuous change of probability measure
and, in particular, makes available Girsanov's theorem.

The original proof in \cite{No_72} is based
on knowing the distribution of the first
exit time of the Wiener process with constant drift
from a shifted positive half-axis.
 Some of other known proofs are even more
  involved (see, for instance, Section 8.1 in \cite{RY_99}).
The latest easier proofs and the history
   revolving around
Novikov's condition are found and thoroughly discussed in \cite{Ru_13}
and \cite{Ru_15}.
Here we   present a completely elementary
 proof of a result that is somewhat
stronger than \eqref{10.17.1}.
\begin{theorem}
                                           \label{theorem 10.16.1}
We have

\begin{equation}
                                           \label{12.9.1}
\nliminf_{\varepsilon\downarrow0}\varepsilon\log 
Ee^{(1-\varepsilon)
\langle M\rangle/2}<\infty\Longrightarrow E\rho=1.
\end{equation}
\end{theorem}

Proof. We start with two known facts (proved at the end
of the paper):
\begin{equation}
                              \label{12.9.2}
 E\rho\leq1;\,\,\,                                                                 
 \exists\varepsilon>0:Ee^{(1+\varepsilon)
\langle M\rangle/2}<\infty\Longrightarrow E\rho=1
\end{equation}
and show a solution of part of Problem 4.3.13
of \cite{Kr_95} following the hint to that problem.
Observe that for small enough $\varepsilon>0$
$$
Ee^{(1+\varepsilon)^{2}\langle (1-\varepsilon)M\rangle/2}
= Ee^{(1-\varepsilon^{2})^{2}\langle M\rangle/2}<\infty,
$$
which by (\ref{12.9.2})   implies that $E\rho((1-\varepsilon)M)=1$.

Then use H\"older's inequality and  
write that for small enough $\varepsilon>0$
and any constant $T\in(0,\infty)$
$$
1=E\rho((1-\varepsilon)M)=Ee^{(1-\varepsilon)
(M-\langle M\rangle/2)}
e^{(1-\varepsilon)\varepsilon\langle
 M\rangle/2}I_{\langle M\rangle\leq T}
$$
$$
+EI_{\langle M\rangle> T}e^{(1-\varepsilon)
(M-\langle M\rangle/2)}
e^{(1-\varepsilon)\varepsilon\langle M\rangle/2}
$$
$$
\leq(E\rho)^{1-\varepsilon}
(Ee^{(1-\varepsilon)\langle M\rangle/2}I_{\langle M\rangle\leq T})^
{\varepsilon}+
(E\rho I_{\langle M\rangle> T})^{1-\varepsilon}
(Ee^{(1-\varepsilon)\langle M\rangle/2})^
{\varepsilon}.
$$
 
As $\varepsilon\downarrow0$, we get $$1
\leq E\rho+{\rm const} \,
E\rho I_{\langle M\rangle> T},$$  which gives $1\leq E\rho$
after letting $T\rightarrow\infty$.  This
together with the first relation in (\ref{12.9.2})
 implies
our statement (\ref{12.9.1}). The theorem is proved.

  Assertion  (\ref{12.9.1})  
 is stronger than \eqref{10.17.1}. 
\begin{example}
                                       \label{example 10.17.1}
 Take a one-dimensional Wiener 
process $w_{t}$ and let $\tau$
be the first exit time of $w_{t}$
from $(-\pi,\pi)$. Take $\lambda\in(0,1/8)$ and observe that
$u_{\lambda}(x)=\cos(\sqrt{2\lambda}x)$
satisfies $(1/2)u''+\lambda u=0$
on $[-\pi,\pi]$.
It follows by It\^o's formula
that $m^{\lambda}_{t}:= u_{\lambda}(w_{t\wedge\tau})
\exp(\lambda(t\wedge\tau))$ is a martingale and since 
$u_{\lambda}$ is bounded away from zero ($\lambda<1/8$),
there is a $\varepsilon>0$ such that, for any $t<\infty$,
$\varepsilon E\exp(\lambda(t\wedge\tau))\leq
Em^{\lambda}_{t}= 1$. As $t\to\infty$, we get
$\varepsilon E\exp \lambda \tau \leq 1$,
which by the dominated convergence theorem
allows us to send $t\to\infty$ in $Em^{\lambda}_{t}= 1$
and obtain that
$$
\cos(\sqrt{2\lambda}\pi)E\exp (\lambda\tau)=1,\quad
E\exp (\lambda\tau)=\big[\cos(\sqrt{2\lambda}\pi)\big]^{-1}
$$
if $\lambda<1/8$ and then $E\exp (\lambda\tau)=\infty$
if $\lambda= 1/8$.
Therefore, for the martingale $M_{t}=w_{t\wedge\tau}/4$
one easily finds that as
$\varepsilon\downarrow0$
$$
[Ee^{(1-\varepsilon)\langle M\rangle/2)}
]^{\varepsilon}=
[Ee^{(1-\varepsilon)\tau/8}]^{\varepsilon}
=\big[\cos \frac{\sqrt{1-\varepsilon}}{2}\pi
\big]^{-\varepsilon}\rightarrow1,
$$
so that the assumption in (\ref{12.9.1}) is 
satisfied, whereas
$E\exp(\langle M\rangle/2)=\infty$ and  
 Novikov's criterion is not
applicable.
\end{example}

\mysection{Refined Novikov's conditions}

In \cite{No_79} Novikov (1979)  relaxes the  conditions
 in \eqref{10.17.1}
and shows that, for any constant $c\geq0$,
\begin{equation}
                                           \label{10.17.2}
Ee^{\langle M\rangle/2-c\langle M\rangle^{1/2}}
<\infty\Longrightarrow 
E\rho=1.
\end{equation}
This condition is applicable  in Example \ref{example 10.17.1},
although it is not very easy to see that. We need to know
the tail of the distribution of $\tau$. On the other hand,
the elementary inequality:
$c\langle M\rangle^{1/2}\leq (\varepsilon/2) 
\langle M\rangle+c^{2}\varepsilon ^{-1}$
implies that
$$
I:=Ee^{\langle M\rangle/2-c\langle M\rangle^{1/2}}
<\infty\Longrightarrow 
\nliminf_{\varepsilon\downarrow0}\varepsilon\log 
Ee^{(1-\varepsilon)
\langle M\rangle/2}\leq \log I-c^{2}<\infty.
$$
 
In the same article \cite{No_79} Novikov gives
a more elaborated condition
\begin{equation}
                                           \label{10.17.3}
Ee^{\langle M\rangle/2-g(\langle M\rangle^{1/2})} 
<\infty\Longrightarrow 
E\rho=1,
\end{equation}
where $g$ belongs to the lower Kolmogorov class,
for instance, 
$$
g(\tau)=\sqrt{2\tau\log\log\tau}
$$
for large $\tau$. This condition is, of course,
much weaker than ours, but checking it, generally,
 requires much more
knowledge about the distribution of $\langle M\rangle$.

On the one hand, \eqref{10.17.3} provides the most general
condition in terms of the distribution
of $\langle M\rangle$. But on the other hand, this
 distribution, generally,  has little to do with the equality
$E\rho=1$. Indeed, if $\tau$ is the first time the Wiener
process $w_{t}$ hits point 1 and $M_{t}=w_{t\wedge\tau}$,
then by Kazamaki's criterion (see \cite{Ka_77})
we have $E\rho=1$ and at the same time even $
E\sqrt{\langle M\rangle}=E\sqrt{\tau}=\infty$
not to mention any exponential moments.
By the way, everything said about \eqref{12.9.1}
 has its natural counterpart
for Kazamaki's criterion (see 
http://arxiv.org/abs/math/0207013).

The conditions described above usually are interesting
not exclusively in their own rights but in connection
with the problem of absolute continuity of
the distribution of a stochastic process with respect to the
Wiener measure. For instance, let $\xi_{t}$ be a (adapted) solution
of a stochastic equation $d\xi_{t}=dw_{t}
+b(\xi_{\cdot},t)\,dt
$, $\xi_{0}=0$, with  
   nonanticipating $b(\xi_{\cdot},t)$ and we are interested
to know when its distribution $\mu_{\xi_{\cdot}}$ on $C(0,1)$
is absolutely continuous with respect to the distribution
$\mu_{w _{\cdot}}$  on $C(0,1)$ of the Wiener process
$  w_{t}$.
  According to Theorem 6 of \cite{LSh_72}, $\mu_{\xi_{\cdot}}\ll
\mu_{w _{\cdot}}$ iff
\begin{equation}
                                               \label{2.1.2}
P\big(\int_{0}^{1}b^{2}(\xi_{\cdot},t)\,dt<\infty\big)=1.
\end{equation}
Under this condition for any nonnegative
measurable function $f(x_{\cdot})$ on $C(0,1)$
\begin{equation}
                                               \label{2.1.1}
Ef(\xi _{\cdot})=Ef(w^{x}_{\cdot})e^{\phi},
\quad \phi=\int_{0}^{1}b(w^{x}_{\cdot},t)\,dw_{t}
-(1/2)\int_{0}^{1}b^{2}(w^{x}_{\cdot},t)\,dt.
\end{equation}
In particular, $Ee^{\phi}=1$. Thus \eqref{2.1.2}
implies that $Ee^{\phi}=1$.
 More general
multidimensional
equations are treated in \cite{LSh_01}.  

In conclusion, for the sake of completeness
we prove \eqref{12.9.2} following 
the proof of Lemma 3 in \cite{LSh_72}.
This shows that unlike \cite{No_72} and \cite{No_79}
 no specific information
about the Wiener process is needed to get our main result.

That $E\rho\leq1$ follows from the fact that $\rho_{t}$
is a local martingale (It\^o's formula) and the fact that
$\rho_{t}\geq0$, so that it is a supermartingale,
has the limit as $t\to\infty$ and Fatou applies.

 Next,
for any $p,r >1$, stopping time $\tau$ such that $\rho_{t\wedge\tau}$
is a martingale, and  
$t\in[0,\infty)$, we easily obtain by the H{\"{o}}lder inequality 
$$ 
E\rho_{t\wedge\tau}^p \leq \bigg(E 
\rho_{t\wedge\tau}(prM)\bigg)^{1/r}
  \bigg(E\exp(\frac{rp-1}{2}p\frac{r}{r-1}
\langle M\rangle_{t\wedge\tau} )\bigg)^
{(r-1)/r}.
$$
By \thetag4, the first factor on the right is at most $1$.
The second factor can be made bounded uniformly with respect to 
$t $, since $\langle M\rangle_{t}\leq\langle M\rangle$
and for $p=1+\delta,\,r=1+\sqrt{\delta}$
and sufficiently small $\delta >0$, it is not difficult to see
that the coefficient of $\langle M\rangle_{t\wedge\tau}$ can be made smaller than
$(1+\varepsilon)/2$. Fix $\delta,p$, with these properties.
Then by Doob's inequality $E\sup_{t\leq\tau}\rho_{t}^{p}\leq N$,
where the constant $N$ is independent of $\tau$. This yields that
$E\sup_{t<\infty}\rho_{t}^{p}\leq N$,
the local martingale $\rho_{t}$ is bounded by a summable
function independent of $t$ and $E\rho=\lim_{n\to\infty}
E\rho_{\tau_{n}}=1$, where $\tau_{n}\to\infty$ is any
localizing sequence for $\rho_{t}$.

{\bf Acknowledgment}. The author is sincerely grateful
to the referee for the comments and suggestions
which certainly helped improve the presentation.

\end{document}